OpenStudy (anonymous):

Find a polynomial with integer coefficients that satisfies the given conditions... P has degree 3 and zeros 2 and i.

OpenStudy (anonymous):

I have this set up... (x-2)(x+i)(x-i) Need help from there...if this is even right lol

OpenStudy (mathmate):

Yes, the polynomial is correct. All you have to do is to expand it to the standard form. Note that (x+i)(x-i)=x^2-i^2.

OpenStudy (anonymous):

So, (x-2)(x^2-i^2)=.... one second

OpenStudy (anonymous):

Well i^2= -1 right?

OpenStudy (mathmate):

yep!

OpenStudy (anonymous):

So it would be (x-2)(x^2-(-1)

OpenStudy (anonymous):

Or (x-2)(x^2+1) ?

OpenStudy (mathmate):

Yep, expand it!

OpenStudy (anonymous):

x^3-2x^2+x-2 ?

OpenStudy (mathmate):

Looks good to me!

OpenStudy (anonymous):

Thanks!

OpenStudy (mathmate):

yw!